OpenStudy (anonymous):

Solve the following equation.Describe solution set as contradiction, identity, or conditional. 3(6-x)+2x=14+4(x-1)-5x

OpenStudy (anonymous):

Contradiction. Expand the left-hand side: 18-3x+2x = 18 - x Now expand the right-hand side: 14 + 4x - 4 - 5x = 10 -x Compare the LHS with RHS: \[18-x=10-x\]add x to both sides and you get\[18=10\]which is a contradiction.